OpenStudy (anonymous):

Factor completely p4 – 16

OpenStudy (turingtest):

Use the difference of squares formula\[x^2-a^2=(x-a)(x+a)\]and note that\[x^4=(x^2)^2\]if you do this correctly you should be able to do it twice.

OpenStudy (anonymous):

i dont getit

OpenStudy (anonymous):

what is a?

OpenStudy (turingtest):

asnaseer is helping you on the other post listen to him...

OpenStudy (lollylau):

(p-2)(p+2)(p^2+4)